DELAYED IONIZATION OF UO_{2} AND PFI-ZEKE SPECTRA FOR UO_{2}^{+}

\underline{JEREMY M. MERRITT}, JIANDE HAN, AND MICHAEL C. HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322.

The decay dynamics of threshold ionization of UO_{2} have been examined using a 1+1 REMPI technique. Ionization is accomplished via resonant excitation of the 2g ← X ^{3}Φ2u transition of UO_{2} at 314.09 nm, followed by ionization just above threshold using a second tunable dye laser operating at λ ≤ 573 nm. While the peak widths for U^{+} and UO^{+} in the time of flight mass-spectrum are limited by the experimental resolution, the peak for UO_{2}^{+} is significantly broader, showing a characteristic decay towards longer flight times, which is the result of delayed ionization. The lifetime of this process is approximately 100 nanoseconds (at threshold) and decreases exponentially as the ionizing laser is tuned higher in energy.

The ionization potential (IP) of UO_{2} has been measured previously, namely 6.128(3) eV, using photoionization efficiency curves. Refinement of the IP using mass analyzed threshold ionization (MATI) spectra was unsuccessful, as the delayed ionization may have interfered with this detection technique [J. Han et al. J. Chem. Phys. 120, 5155 (2004)]. Here we report high-resolution pulsed-field ionization zero kinetic energy photoelectron (PFI-ZEKE) spectra which have allowed a more accurate determination of the IP of UO_{2}, as well as providing a spectroscopic probe of UO_{2}^{+}. Accurate values for the electronic term energies and vibrational frequencies of UO_{2}^{+} have been obtained. A comparison of the current experimental results with theoretical calculations will be presented.